a) Calculate the energy required to excite a particle in a cubic box of edge length 1A^° from its ground state to its first excited state. b) what us the reason for the increase in the spectral frequency as be increases?
1
Expert's answer
2020-10-30T06:42:35-0400

a) Potential well of width (L) = 1 × 10–10 m

E=b^2/8mL^2 = (6.63x10-34)^2/8x9.1x10-31 x 10-10 = 0.6038 x 10-17

b) The Doppler effect (or the Doppler shift) is the change in frequency of a wave in relation to an observer who is moving relative to the wave source.